A bar 100 cm long, with insulated sides, has its ends kept at 0
1
Expert's answer
2021-07-08T11:50:51-0400

A bar 100 cm long, with insulated sides, has its ends kept at 0° Celsius and 100° Celsius until

steady state condition prevails. The two ends are then suddenly insulated and kept so. Find

the temperature distribution.​

Solution

The heat flow flow equation is

"\nabla ^2u(x,t)={1\over \alpha ^2}{\partial u\over \partial t}...................(i)"

where "u(x, t)" is the temperature. Because the sides of the bar are insulated, the heat flows only in the "x" direction; the same happens for a slab of finite thickness but infinitely large. The initial condition is "u(x, 0) = 100" and the boundary condition for "t > 0" is "u(0, t) = u(100, t) = 0" . We search for a solution of the form "u(x, t) = F(x)T(t)" ; the differential equation "(i)" gives

"{∂ ^2F\over ∂x^2} + k ^2F = 0 , \implies F = A cos kx + B sin kx.............(ii)"

"{∂T\over ∂t} = −k^2α^2T , \implies T = e^{−k^2α^2t}....................(iii)"

Because of the boundary condition in "x = 0" , we have "A = 0" . To satisfy "u(100, t) = 0" , we find discrete values of "k:"

"sin 100k = 0 ,\implies k_n ={nπ\over100}," ".........(iv)" ,for "n = 1, 2, 3, . . ."

Finally, we must satisfy the initial condition

"u(x, 0) =\displaystyle\sum_{n=1}^∞b_n sin{nπx\over 100}= 100 ...............(v)" ,for "0 \u2264 x \u2264 10."

Now,solving for "b_n" we have

"b_n ={2\over 100} \int_0^{100 }sin{nπx\over 100}dx = −{200\over nπ}cos{nπx\over 100}\mid_0^{100}"

"={200\over nπ}[1 − (−1)^n]"

"= \begin{cases}
 {400\over \pi n}&\text{for }{odd} {n} \\
 0 &\text{for } {even} n
\end{cases}" "..............(vii)"

after using "cos (n\u03c0 )= (\u22121)^n" for integer "n" . The solution to the differential equation "(i)" with the given boundary conditions is then


"u (x,t)={400\over π}\sum_ {oddn}{1\over n}e^{−({nπα\over 100})^2t }sin {nπx\over 100}"


where “odd n” means a sum over n = 1, 3, 5, . . ..